# USTM: Unified Spatial and Temporal Modeling for Continuous Sign Language Recognition

### v1

@@ II Related Work
CSLR Approaches. Single-stream CSLR approaches process only RGB videos through a single pathway, aiming to capture spatial-temporal dependencies without leveraging additional modalities or parallel streams. Cheng et al. [cheng2020fully] proposed a fully convolutional framework with gloss feature enhancement to improve alignment. To better capture local motion, Xie et al. [xie2023multi] introduced a multi-scale temporal similarity fusion network using position-aware convolutions. VAC loss [min2021visual] enforces alignment consistency between modules, improving feature quality. Hu et al. [hu2023self] emphasized hand/face regions using spatial-temporal attention, while Hao et al. [hao2021self] used mutual distillation to reduce CTC spikes. Temporal Lift Pooling (TLP) [hu2022temporal] improves temporal downsampling, and CorrNet [hu2023continuous] models cross-frame motion via dynamic correlations. To handle long sequences, Cui et al. [cui2023spatial] proposed a spatial-temporal transformer using frame chunking. Wang et al. [wang2025continuous] proposed STNet, which strengthens frame-wise and multi-scale spatio-temporal representations through a spatial resonance module that captures global common spatial features across adjacent frames, together with a multi-temporal perception module. Similarly, Gan et al. [gan2024signgraph] proposed SignGraph, where sign sequences are represented as graphs to capture intra- and inter-frame cross-region interactions across multiple scales. Most recently, Alyami et al. [alyami2025swin] enhanced Swin Transformers with multiscale temporal modules for fine-grained motion modeling.
Multi-modal CSLR models leverage additional cues such as pose, optical flow, and hand/face features to enhance recognition performance. Cui et al. [cui2019deep] combined CNN-RNN architectures with iterative RGB-flow fusion to better capture temporal dependencies. SignBERT [zhou2021signbert] integrates hand crops with transformer-based models to improve generalization across different signers. Zhou et al. [zhou2021spatial] proposed a spatial-temporal multi-cue network that jointly models multiple visual cues through dedicated spatial and temporal modules. C2SLR [zuo2022c2slr] employs keypoint-guided spatial attention and sentence embedding consistency constraints to enhance both visual and sequential modules. TwoStream-SLR [chen2022two] and CVT-SLR [zheng2023cvt] explore dual-stream and contrastive visual-textual alignment strategies, combining raw RGB, keypoint sequences, and pretrained language modules for more robust cross-modal representations. Multi-channel transformers [camgoz2020multi] explicitly model intra- and inter-channel relationships of different articulators, while temporal deformable convolutional sequence learning [papadimitriou2020multimodal] fuses multiple spatio-temporal streams for end-to-end recognition. Finally, frameworks such as SAM-SLR [jiang2021skeleton] incorporate graph-based skeleton modeling alongside RGB and depth cues.
Multi-stream models are designed with multiple parallel pathways to process the same type of input modality, enabling the extraction of complementary features. In the context of CSLR, these models have demonstrated considerable improvements in capturing both spatial and temporal information. For instance, the SlowFast network [ahn2024slowfast] employs a dual-pathway architecture, where one pathway operates at a high temporal resolution to capture dynamic motion, while the other focuses on spatial details, such as hand shapes and facial expressions. AdaBrowse [hu2023adabrowse] introduces a lightweight adaptive mechanism that selectively processes the most informative temporal subsequences and dynamically adjusts input resolution. Similarly, cross-modal augmentation strategies [pu2020boosting] generate pseudo video-text pairs to enhance alignment between video features and textual labels. Beyond RGB inputs, keypoint-based multi-stream models, such as MSKA-SLR [guan2405multi], leverage skeletal information to reduce background noise sensitivity and computational complexity. More recently, a graph-based multi-stream approach, namely MSD-GCN [rastgoo2025multi], integrated adaptive motion-aware attention, signal-enhanced feature extraction, and self-supervised pretraining to capture fine-grained motion patterns and long-term dependencies.
Figure 2: Architecture of TAPE adapter.
Transformer Architectures for Video Understanding. Vision Transformers (ViTs) [dosovitskiy2020image] introduced a new paradigm in visual learning by applying self-attention over image patches to effectively capture long-range dependencies without convolution. While ViT showed strong performance on large-scale datasets, follow-up models, such as DeiT [touvron2021training], T2T-ViT [yuan2021tokens], and PVT [wang2021pyramid], improved performance in data-constrained settings via distillation, better tokenization, and hierarchical design. Later variants like Swin Transformer [liu2021swin] and CvT [wu2021cvt] introduced local inductive biases and convolutional priors, enabling ViTs to close the performance gap with CNNs on dense prediction tasks. These developments paved the way for transformer-based architectures in video understanding, where modeling both spatial and temporal dynamics is essential.
To extend ViTs to video, several models have been proposed to explicitly capture temporal dependencies. TimeSformer [bertasius2021space] factorized attention into spatial and temporal components, demonstrating competitive performance on action recognition tasks. ViViT [arnab2021vivit] examined various attention schemes, including joint and factorized attention with tubelet embeddings. MViT [fan2021multiscale] introduced multi-scale attention and pooling mechanisms, while Uniformer [li2023uniformer] unified convolution and self-attention for efficient spatial-temporal modeling. More recently, STAdapter [pan2022st] proposed injecting temporal reasoning into spatial backbones using lightweight temporal adapters. This approach offers a modular and parameter-efficient solution for adapting pretrained vision transformers to video-based tasks.
@@ IV-B Comparison with Other Works
RGB-Based Frameworks. Table I compares the performance of our USTM framework with prior single-stream RGB-based methods. As shown in the table, USTM variants demonstrate consistent improvements across all datasets. On Phoenix14, USTMtiny achieves WERs of 18.1 and 18.3 on the dev and test sets, respectively. USTMsmall achieves the same scores, while USTMbase improves to 17.4 on dev and 17.7 on test. Our smallest model remains highly competitive, trailing CorrNet+ by only 0.1 on both dev and test. USTMbase surpasses the previous SOTA CorrNet+ with WER of 18.0 on dev and 18.2 on test, setting a new SOTA on this dataset. Notably, USTMtiny, which uses the tiny Swin transformer backbone, outperforms Swin-MSTP that relies on the larger Swinsmall variant which underscores the efficacy of the proposed approach.
@@
Method Phoenix14 Phoenix14T CSL-Daily
@@
FCN [cheng2020fully] 23.7 23.9 23.3 25.1
mLTSF-Net [xie2023multi] 22.9 23.0
VAC [min2021visual] 21.2 22.3
MSTNet [xie2023multi] 20.3 21.4
SEN [hu2023self] 19.5 21.0 19.3 20.7
SMKD [hao2021self] 20.8 21.0 20.8 22.4
TLP [hu2022temporal] 19.7 20.8 19.4 21.2
SSSLR [jang2023self] 20.9 20.7 20.5 22.3
CTCA [guo2023distilling] 19.5 20.3 19.3 20.3 31.3 29.4
ST-Transformer [cui2023spatial] 19.9 19.9
CorrNet [hu2023continuous] 18.8 19.4 18.9 20.5 30.6 30.1
CorrNet+ [hu2024corrnet+] 18.0 18.2 17.2 19.1 28.6 28.2
Swin-MSTP [alyami2025swin] 18.1 18.7 18.7 19.7 28.3 27.1
CLIP-SLA [alyami2025swin] 19.7 19.3 19.8 19.4 26.0 25.8
USTMtiny (Ours) 18.1 18.3 17.8 19.4 27.8 27.6
USTMsmall (Ours) 18.1 18.3 17.4 19.3 27.7 26.4
USTMbase (Ours) 17.4 17.7 17.6 18.9 27.7 26.5
@@
Multi-Stream Frameworks. Table II compares the performance of our USTM framework with prior multi-stream frameworks that use multiple streams to capture varied feature representations. On Phoenix14, USTMbase surpasses all previous methods, including SlowFastSign, which achieved WERs of 18.0 and 18.3 on the dev and test sets, respectively. Our smaller variants achieve performance on par with SlowFastSign, demonstrating the efficiency of our single-stream design. On Phoenix14T, our models surpass previous methods on the dev split, while trailing the prior best on the test split by a WER of 0.2. On CSL-Daily, USTMsmall trails the previous best method by a margin of 2.2 on the dev split and 1.5 on the test split. The performance gap between our models and SlowFastSign can be attributed to two main factors. First, SlowFastSign is pretrained on the large-scale Kinetics-400 dataset, which provides a strong spatiotemporal prior, while our models are only initialized with ImageNet weights. Second, the SlowFast architecture is explicitly designed to model temporal dynamics at multiple scales throughout the network, which inherently benefits continuous sign language recognition. Despite these differences, our single-stream architecture remains highly competitive and demonstrates strong generalization across datasets without relying on multi-stream inputs or specialized temporal pretraining.
@@
Method Phoenix14 Phoenix14T CSL-Daily
@@
CMA [pu2020boosting] 23.9 24.0 24.1 24.3
AdaBrowse [hu2023adabrowse] 19.6 20.7 19.5 20.6 31.2 30.7
MSKA [ahn2024slowfast] 21.7 22.1 20.1 20.5 28.2 27.8
SlowFastSign [ahn2024slowfast] 18.0 18.3 17.7 18.7 25.5 24.9
USTMtiny (Ours) 18.1 18.3 17.8 19.4 27.8 27.6
USTMsmall (Ours) 18.1 18.3 17.4 19.3 27.7 26.4
USTMbase (Ours) 17.4 17.7 17.6 18.9 27.7 26.5
Multi-Modal Frameworks. We also compare the performance of the USTM framework with other methods that utilize more than one modality. Table III compares the results of our framework with prior multi-modal methods. As shown in the table, most of the multi-modal methods use combine RGB with pose information. On the Phoenix14 dataset, USTMbase achieves the best results, surpassing all previous works. Notably, even our smallest variant, USTMtiny, outperforms TwoStream-SLR the SOTA multi-modal CSLR, achieving 18.1 in dev and 18.3 in test. USTMsmall also performs competitively with the same scores, matching or slightly improving over prior SOTA. For Phoenix14T, USTM improves over TwoStream-SLR with 17.7 in dev and 19.3 in test. On CSL-Daily, USTMsmall attains the lowest WER, surpassing TwoStream-SLR, with 33.1 in dev and 32.0 in test sets.
@@
Method Modality Phoenix14 Phoenix14T CSL-Daily
@@
DNF [cui2019deep] RGB+Optical Flow 23.1 22.8 32.8 32.4
SignBERT [zhou2021signbert] RGB+Pose 20.1 20.2 33.6 33.1
STMC [zhou2021spatial] RGB+Pose 20.6 21.5 19.6 21.0
C2SLR [zuo2022c2slr] RGB+Pose 20.5 20.4 20.2 20.4
MSTCSLRN [li2022multi] RGB+Gloss 22.8
CVT-SLR [zheng2023cvt] RGB+Pose 19.8 20.1 19.4 20.3
TwoStream-SLR [chen2022two] RGB+Pose 18.4 18.8 17.7 19.3 33.1 32.0
USTMtiny (Ours) RGB 18.1 18.3 17.8 19.4 27.8 27.6
USTMsmall (Ours) RGB 18.1 18.3 17.4 19.3 27.7 26.4
USTMbase (Ours) RGB 17.4 17.7 17.6 18.9 27.7 26.5

### v2

@@ II Related Work
Late Stage Temporal Modeling. Most existing CSLR approaches address the recognition problem by first extracting spatial representations independently from individual frames and subsequently performing temporal sequence modeling on the resulting frame-level features. We refer to this paradigm as late stage temporal modeling. In this setting, temporal dependencies are modeled on top of compressed spatial representations using sequential modules such as temporal convolutions, recurrent networks, attention mechanisms, or transformers, allowing models to leverage strong image encoders while deferring temporal reasoning to later stages. Within this paradigm, spatial encoding is typically performed using off-the-shelf image backbones. The most commonly adopted backbone is ResNet-18, which serves as the primary spatial encoder in a large portion of late stage temporal modeling approaches [zhou2021signbert, zheng2023cvt, min2021visual, hu2022temporal, wang2025continuous, yang2025acmc]. Deeper ResNet variants, such as ResNet-34, are also employed in some works to enhance spatial representation capacity [zhu2025continuous]. Lightweight architectures like MobileNetV2 are explored to reduce computational cost [ranjbar2025continuous], while more recent studies adopt transformer-based image encoders, including MViT [nam2025importance] and Swin Transformers [alyami2025swin], to capture hierarchical and multi-scale spatial cues.
Early work adapts image-centric backbones by augmenting them with lightweight temporal mechanisms. Cheng et al. [cheng2020fully] introduced a fully convolutional framework with gloss feature enhancement to improve alignment. Cui et al. [cui2019deep] proposed an iterative CNN-RNN training strategy to better capture temporal dependencies. Zhou et al. [zhou2021signbert] presented a self-emphasizing network that integrates hand-centric cues to improve robustness across signers. Spatial and temporal cues were jointly modeled through multi-cue designs in Zhou et al. [zhou2021spatial]. Several methods focused on improving temporal aggregation without departing from image backbones. These include multi-scale local temporal similarity fusion [xie2023multi], visual alignment constraints [min2021visual], self-mutual distillation to reduce CTC spikes [hao2021self], and Temporal Lift Pooling for improved temporal downsampling [hu2022temporal]. Cui et al. [cui2023spatial] introduced a spatial-temporal transformer that operates on frame chunks. Other extensions explored consistency constraints and auxiliary supervision, such as C2SLR [zuo2022c2slr], multi-channel transformers for multi articulatory modeling [camgoz2020multi], temporal deformable convolutional fusion of multimodal streams [papadimitriou2020multimodal], adaptive temporal browsing in AdaBrowse [hu2023adabrowse], and cross-modality augmentation to improve video-text alignment [pu2020boosting]. Contrastive visual-textual learning with variational alignment was further explored in CVT-SLR [zheng2023cvt].
Recent methods strengthen temporal reasoning within modern image transformers and explore hybrid temporal-spatial modules. Alyami et al. [alyami2025swin] proposed Swin-MSTP, which augments Swin Transformers with multi-scale temporal perception modules for fine-grained motion modeling. Wang et al. [wang2025continuous] introduced a multi-scale spatial-temporal feature enhancement framework that reinforces frame-wise representations and cross-frame interactions. Huang et al. [huang2025dual] proposed a dual-stage temporal perception network combining multi-scale local temporal modules with global-local temporal relational modules for richer video features. Tran et al. [tran2025generalizable] presented a pose-only CSLR framework leveraging local temporal convolutions and region-aware pose encoding for robust generalization. Nam et al. [nam2025importance] systematically investigated the impact of different facial regions on recognition, showing the mouth is the most critical non-manual feature. Ranjbar et al. [ranjbar2025continuous] introduced intra-inter gloss attention, exploiting relationships within and between glosses for improved transformer-based CSLR. Yang et al. [yang2025acmc] proposed ACMC, an adaptive cross-modal multi-grained contrastive learning framework for fine-grained visual-text alignment. Finally, Zhu et al. [zhu2025continuous] introduced a motor attention mechanism combined with frame-level self-distillation to capture dynamic motion changes, enhancing feature representation and inference robustness.
Early Stage Temporal Modeling. In contrast to late stage temporal modeling, early stage temporal modeling refers to CSLR approaches that integrate temporal reasoning directly into the feature extraction process, rather than applying it solely on top of precomputed frame-level representations. In this paradigm, spatial and temporal cues are jointly modeled from the early layers of the network, enabling motion patterns, inter-frame dependencies, and articulatory dynamics to influence feature learning throughout the encoding pipeline. Methods following this approach often rely on spatio-temporal backbones or architectures that explicitly encode temporal information during spatial feature extraction. Representative designs include 3D convolutional networks such as ResNet2+1D-18 [jiang2021skeleton], S3D [chen2022two], and SlowFast-101 [ahn2024slowfast], which jointly model appearance and motion across frames. Other approaches incorporate temporal interactions through alternative formulations, such as spatio-temporal graph neural networks for structured sign representation [gan2024signgraph], attention-based multi-stream keypoint encoders [guan2405multi], or by injecting temporal adapters within 2D CNN backbones to enable early temporal coupling [hu2023continuous].
Jiang et al. [jiang2021skeleton] introduced a skeleton-aware multi-modal framework that leverages graph-based skeleton cues alongside RGB and depth inputs. Chen et al. [chen2022two] proposed TwoStream-SLR, a dual-pathway network that separately encodes raw video frames and keypoint sequences, and uses bidirectional connections, sign pyramid networks, and frame-level self-distillation for enhanced spatio-temporal learning. Hu et al. [hu2023continuous] presented CorrNet, which dynamically computes correlation maps between adjacent frames to emphasize body trajectories, though the approach may not fully capture long-range temporal dependencies. Gan et al. [gan2024signgraph] proposed SignGraph, representing sign sequences as graphs to model intra-frame and inter-frame cross-region interactions. Similarly, Ahn et al. [ahn2024slowfast] developed a SlowFast network with dual pathways for high-temporal and high-spatial resolution processing. Multi-stream keypoint attention models [guan2405multi] and diffusion graph convolution networks with adaptive motion-aware attention [rastgoo2025multi] further explore the combination of skeletal cues, multi-stream fusion, and self-supervised pretraining to capture fine-grained motion and long-term dependencies. Zhou et al. [zhou2025scaling] introduced large-scale multimodal pretraining, integrating pose and text cues to enhance visual representation learning and contextual understanding.
These approaches embed temporal reasoning directly within early feature extraction via 3D convolutions, temporal adapters, spatio-temporal graphs, or attention mechanisms, enabling motion-aware representations from the outset. Despite this, early integration of temporal modeling in CSLR remains relatively underexplored.
@@ IV-B Comparison with Other Works
RGB-Based Frameworks. Table I compares the performance of our USTM framework with prior single-stream RGB-based methods. As shown in the table, USTM variants demonstrate consistent improvements across all datasets. On Phoenix14, USTMtiny achieves WERs of 18.1 and 18.3 on the dev and test sets, respectively. USTMsmall achieves 17.9 on dev and 17.6 on test, while USTMbase improves to 17.4 on dev and 17.7 on test. Our smallest model remains highly competitive, trailing CorrNet+ by only 0.1 on both dev and test. USTMsmall surpasses the previous SOTA CorrNet+ with WER of 18.0 on dev and 18.2 on test, setting a new SOTA on this dataset. Notably, USTMtiny, which uses the tiny Swin transformer backbone, outperforms Swin-MSTP that relies on the larger Swinsmall variant which underscores the efficacy of the proposed approach.
@@
Method Venue Phoenix14 Phoenix14T CSL-Daily
@@
FCN [cheng2020fully] ECCV 2020 23.7 23.9 23.3 25.1
mLTSF-Net [xie2023multi] PR 2023 22.9 23.0
VAC [min2021visual] ICCV 2021 21.2 22.3
MSTNet [xie2023multi] PR 2023 20.3 21.4
SEN [hu2023self] AAAI 2023 19.5 21.0 19.3 20.7
SMKD [hao2021self] ICCV 2021 20.8 21.0 20.8 22.4
TLP [hu2022temporal] ECCV 2022 19.7 20.8 19.4 21.2
SSSLR [jang2023self] ICASSP 2023 20.9 20.7 20.5 22.3
CTCA [guo2023distilling] CVPR 2023 19.5 20.3 19.3 20.3 31.3 29.4
ST-Transformer [cui2023spatial] CIS 2023 19.9 19.9
CorrNet [hu2023continuous] CVPR 2023 18.8 19.4 18.9 20.5 30.6 30.1
CorrNet+ [hu2024corrnet+] arXiv 2024 18.0 18.2 17.2 19.1 28.6 28.2
Swin-MSTP [alyami2025swin] Neurocomp. 2025 18.1 18.7 18.7 19.7 28.3 27.1
CLIP-SLA [alyami2025clip] CVPR 2025 19.7 19.3 19.8 19.4 26.0 25.8
MP-SLR[zhou2025scaling] TPAMI 2025 21.2 21.2 20.1 21.3 28.6 27.9
STNet[wang2025continuous] IEEE Acc. 2025 19.2 19.3 19.2 19.8 28.3 27.2
IIGA[ranjbar2025continuous] MMTA 2025 20.1 20.4 - - - -
ACMC [yang2025acmc] IMAVIS 2025 17.8 18.4 17.6 18.8 28.3 27.5
USTMtiny (Ours) - 18.1 18.3 17.8 19.4 27.8 27.6
USTMsmall (Ours) - 17.9 17.6 17.4 19.3 27.7 26.4
USTMbase (Ours) - 17.4 17.7 17.6 18.9 27.7 26.5
@@
Multi-Stream Frameworks. Table II compares the performance of our USTM framework with prior multi-stream frameworks that use multiple streams to capture varied feature representations. On Phoenix14, USTMsmall surpasses all previous methods, including SlowFastSign, which achieved WERs of 18.0 and 18.3 on the dev and test sets, respectively. Our smaller variants achieve performance on par with SlowFastSign, demonstrating the efficiency of our single-stream design. On Phoenix14T, our models surpass previous methods on the dev split, while trailing the prior best on the test split by a WER of 0.2. On CSL-Daily, USTMsmall trails the previous best method by a margin of 2.2 on the dev split and 1.5 on the test split. The performance gap between our models and SlowFastSign can be attributed to two main factors. First, SlowFastSign is pretrained on the large-scale Kinetics-400 dataset, which provides a strong spatiotemporal prior, while our models are only initialized with ImageNet weights. Second, the SlowFast architecture is explicitly designed to model temporal dynamics at multiple scales throughout the network, which inherently benefits continuous sign language recognition. Despite these differences, our single-stream architecture remains highly competitive and demonstrates strong generalization across datasets without relying on multi-stream inputs or specialized temporal pretraining.
@@
Method Venue Phoenix14 Phoenix14T CSL-Daily
@@
CMA [pu2020boosting] ACM MM 2020 23.9 24.0 24.1 24.3
AdaBrowse [hu2023adabrowse] ACM MM 2023 19.6 20.7 19.5 20.6 31.2 30.7
MSKA [guan2025mska] PR 2025 21.7 22.1 20.1 20.5 28.2 27.8
SlowFastSign [ahn2024slowfast] ICASSP 2024 18.0 18.3 17.7 18.7 25.5 24.9
SlowFastSign (Reproduced) [ahn2024slowfast] ICASSP 2024 18.1 18.4 17.9 18.9 26.9 26.7
USTMtiny (Ours) - 18.1 18.3 17.8 19.4 27.8 27.6
USTMsmall (Ours) - 17.9 17.6 17.4 19.3 27.7 26.4
USTMbase (Ours) - 17.4 17.7 17.6 18.9 27.7 26.5
Multi-Modal Frameworks. We also compare the performance of the USTM framework with other methods that utilize more than one modality. Table III compares the results of our framework with prior multi-modal methods. As shown in the table, most of the multi-modal methods use combine RGB with pose information. On the Phoenix14 dataset, USTMsmall achieves the best results, surpassing all previous works. Notably, even our smallest variant, USTMtiny, outperforms TwoStream-SLR the SOTA multi-modal CSLR, achieving 18.1 in dev and 18.3 in test. USTMsmall also performs competitively with the same scores, matching or slightly improving over prior SOTA. For Phoenix14T, USTM improves over TwoStream-SLR with 17.7 in dev and 19.3 in test. On CSL-Daily, USTMsmall attains the lowest WER, surpassing TwoStream-SLR, with 33.1 in dev and 32.0 in test sets.
@@
Method Venue Modality Phoenix14 Phoenix14T CSL-Daily
@@
DNF [cui2019deep] IEEE ToM 2019 RGB+Op.Flow 23.1 22.8 32.8 32.4
SignBERT [zhou2021signbert] IEEE Acc. 2021 RGB+Pose 20.1 20.2 33.6 33.1
STMC [zhou2021spatial] IEEE ToM 2021 RGB+Pose 20.6 21.5 19.6 21.0
C2SLR [zuo2022c2slr] CVPR 2022 RGB+Pose 20.5 20.4 20.2 20.4
MSTCSLRN [li2022multi] arXiv 2022 RGB+Gloss 22.8
CVT-SLR [zheng2023cvt] CVPR 2023 RGB+Pose 19.8 20.1 19.4 20.3
TwoStream-SLR [chen2022two] NeurIPS 2022 RGB+Pose 18.4 18.8 17.7 19.3 33.1 32.0
USTMtiny (Ours) - RGB 18.1 18.3 17.8 19.4 27.8 27.6
USTMsmall - RGB 17.9 17.6 17.4 19.3 27.7 26.4
USTMbase (Ours) - RGB 17.4 17.7 17.6 18.9 27.7 26.5